OpenStudy (anonymous):

expand (x+5) raise to 4

OpenStudy (campbell_st):

use the coefficients from Pascals triangle 1, 4, 6, 5, 4, 1 then its \[1 \times x^4 + 4 \times x^3 \times 5 + 6 \times x^2 \times 5^2 + 4 \times x \times5^3 + 5^4\] simplifying gives \[(x + 5)^4 = x^4 + 20x^3 + 150x^2 + 500x + 625\] or you could use binomial expansion \[^{n}C _{r} x^{n - r} 5^r\]